OpenStudy (anonymous):

how do u intergrate this? x/(1+x^2)

OpenStudy (anonymous):

show all steps please

OpenStudy (anonymous):

I'll give you a clue. Take : 1+ x^2 = t.

OpenStudy (anonymous):

1+x^2 =t 2x dx/dt = 1 2xdx = dt so now ur problem becomes Integration of dt / 2t (Note : 2 is multiplied and divided to get this form) Now after integrating we get 1/2 log t substitute t = x^2 1/2 log x^2 = log x since it is not definite integral answer is log x +C